### v1

# Making Teams and Influencing Agents: Efficiently Coordinating Decision Trees for Interpretable Multi-Agent Reinforcement Learning

Abstract
Poor interpretability hinders the practical applicability of multi-agent reinforcement learning (MARL) policies. Deploying interpretable surrogates of uninterpretable policies enhances the safety and verifiability of MARL for real-world applications. However, if these surrogates are to interact directly with the environment within human supervisory frameworks, they must be both performant and computationally efficient. Prior work on interpretable MARL has either sacrificed performance for computational efficiency or computational efficiency for performance. To address this issue, we propose HYDRAVIPER, a decision tree-based interpretable MARL algorithm. HYDRAVIPER coordinates training between agents based on expected team performance, and adaptively allocates budgets for environment interaction to improve computational efficiency. Experiments on standard benchmark environments for multi-agent coordination and traffic signal control show that HYDRAVIPER matches the performance of state-of-the-art methods using a fraction of the runtime, and that it maintains a Pareto frontier of performance for different interaction budgets.
1 Introduction
Over the past decade, multi-agent reinforcement learning (MARL) algorithms have achieved state-of-the-art performance in various challenging board and video games (Silver et al. 2016; Vinyals et al. 2019). They have also found success in other sequential decision-making scenarios based on critical real-world domains, including robotics (Orr and Dutta 2023), cybersecurity (Panfili et al. 2018), and traffic signal control (TSC) (Chen et al. 2020). However, the real-world applicability of these algorithms is hampered by two key challenges. First, the deep neural network (NN) architectures that are needed to achieve good performance have thousands to millions of parameters. Second, the behavior of RL agents is difficult to predict and verify due to its dependency on complex state spaces and long time horizons. Thus, human stakeholders understand and trust RL agents less than their simpler counterparts, even if RL yields superior performance (Siu et al. 2021). Conversely, more interpretable representations of policies can help stakeholders to build appropriate levels of trust in RL agents (Druce, Harradon, and Tittle 2021; Zhang, Liao, and Bellamy 2020).
In applications where the safety and verifiability of RL policies is critical (Gilbert et al. 2023; Jayawardana, Landler, and Wu 2021), such as TSC, users may deploy interpretable surrogate policies in place of expert NN policies. Such surrogate policies should be performant — capable of achieving high returns. In MARL, coordinating the training of surrogates is critical for performance: if multiple surrogates are deployed simultaneously, they cannot assume that they are interacting with performant experts, as their performance may be influenced by suboptimal behaviour from others.
At the same time, surrogate policies should be computationally efficient — it should be possible to generate them with minimal environment interactions and runtime. Frequent real-world environment interactions are impractical in domains such as robotics (Finn et al. 2017) and TSC (Zang et al. 2020), and simulators remain computationally intensive if they need to be of sufficient fidelity to guarantee real-world transferrability (Chen et al. 2023). In human-in-the-loop frameworks where users provide oversight to correct undesirable policy behaviour (Mandel et al. 2017), the ability to iterate on surrogate policies in a rapid fashion is also critical (Wu et al. 2023). The desiderata of performance and computational efficiency exist in tension: more complex models capable of stronger performance and coordination capabilities are less efficient (Milani et al. 2024).
Decision trees (DTs) have been considered an attractive model structure for interpretable single-agent RL due to their comprehensibility (Silva et al. 2020). They lie at the core of the imitation learning framework VIPER (Bastani, Pu, and Solar-Lezama 2018), which has been applied to distill NN-based RL policies into DTs in domains such as TSC (Jayawardana, Landler, and Wu 2021; Zhu, Yin, and Chen 2022), autonomous vehicles (Schmidt et al. 2021), and robotics (Roth et al. 2023). However, generalizing VIPER to the MARL setting is challenging. Prior work (Milani et al. 2022) has introduced two multi-agent VIPER algorithms, IVIPER and MAVIPER. IVIPER fails to coordinate the training of DTs, thus sacrificing performance; meanwhile, MAVIPER trains DTs in a coordinated but computationally inefficient manner. This makes DT surrogates generated by these algorithms impractical for deployment.
To this end, we introduce HYDRAVIPER, an efficient method to extract coordinated DT policies for cooperative MARL. Our method makes three key algorithmic contributions: (1) HYDRAVIPER coordinates agent training by jointly resampling the training dataset for each team of cooperative agents. (2) When interacting with the environment to collect a training dataset, HYDRAVIPER adaptively collects critical trajectories closer to convergence. (3) When interacting with the environment for evaluation, HYDRAVIPER uses a multi-armed bandit-based evaluation strategy to identify promising sets of trained surrogates. Experiments on various environments demonstrate that HYDRAVIPER achieves our goal of balancing performance and computational efficiency. HYDRAVIPER also improves the practical applicability of DT-based interpretable MARL policies: users can exchange training time for performance by altering the algorithm’s environment interaction budgets, but its performance remains optimal at different budget levels. Lastly, HYDRAVIPER’s efficiency on large environments can be improved while maintaining coordination by dividing the agent set into mutually influential teams.
2 Related Work
Interpretable Multi-Agent Learning
Past methods for interpretable MARL have focused on using feature importance measures to construct saliency maps (Iqbal and Sha 2019; Heuillet, Couthouis, and Díaz-Rodríguez 2022; Liu, Zhu, and Chen 2023; Motokawa and Sugawara 2023), logical structures (Kazhdan, Shams, and Lio 2020; Wang et al. 2021; Ji, Li, and Xiao 2023), and domain concepts (Zabounidis et al. 2023). Each of these categories of methods has limitations. Feature importances and saliency maps are visua lly clear, but only highlight aspects of the state space without showing how policies use them. Policies based on logical rules and concepts allow users to align the execution of these policies with domain knowledge, but they require extensive feature engineering. By contrast, we learn simple policy representations grounded directly in the environment feature space.
Decision Trees for Reinforcement Learning
Relative to deep NNs, shallow DT policy representations are intrinsically (Molnar 2019) and empirically (Silva et al. 2020) more comprehensible. One line of work in DT-based RL directly trains DT policies (Silva et al. 2020; Topin et al. 2020; Crespi et al. 2023; Liu et al. 2024) using relaxations amenable to direct optimization. However, these methods suffer from training instability and performance degradation. Another line of work follows the VIPER framework (Bastani, Pu, and Solar-Lezama 2018), in which a surrogate DT is trained through imitation learning of a performant expert. Although VIPER has achieved success in single-agent settings (Schmidt et al. 2021; Roth et al. 2023; Jayawardana, Landler, and Wu 2021; Zhu, Yin, and Chen 2022), only two VIPER-based algorithms exist for the multi-agent setting: IVIPER and MAVIPER (Milani et al. 2022). IVIPER independently trains DTs for each agent in a decentralised manner, enjoying computational efficiency at the cost of performance due to lack of coordination. MAVIPER jointly trains DTs in a centralised manner to achieve coordination, but it suffers from computational inefficiency. Neither algorithm balances performance and computational efficiency.
3 Background
Figure 1: (L) Imitation learning in traffic signal control, where a decision tree must be learned to imitate the RL-based signal control policy of each intersection’s controller agent (the seven-intersection Ingolstadt corridor TSC environment from Section 5.1 is shown, with intersections numbered). (R) The HYDRAVIPER framework, in which DT students are trained independently using a jointly resampled dataset of trajectories collected from an environment and relabeled by an NN expert.
Markov Games
We model agent cooperation as a team-based Markov game. A Markov game for $N$ agents consists of a set of states $\mathcal{S}$ with initial state distribution $\rho:\mathcal{S}\rightarrow[0,1]$, and a set of actions $\mathcal{A}_{1},...,\mathcal{A}_{N}$ and observations $\mathcal{O}_{1},...,\mathcal{O}_{N}$ (consisting of features correlated with the state) for each agent $i$. The agent set is partitioned into disjoint teams $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\subseteq[N]$. Each agent chooses actions according to a policy $\pi_{i}:\mathcal{O}_{i}\rightarrow\mathcal{A}_{i}$. After agents simultaneously execute actions $\mathbf{a}$, the environment produces the next state based on the state transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\ldots\times\mathcal{A}_{N}\rightarrow\mathcal{S}$, a joint reward for each team $\mathcal{T}_{\ell}$ based on $R_{\ell}:\mathcal{S}\times\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\rightarrow\mathbb{R}$, and private observations $O_{i}:\mathcal{S}\rightarrow\mathcal{O}_{i}$. Each agent maximises its team’s return $\mathcal{R}_{\ell}=\sum_{t=0}^{T}\gamma^{t}R_{\ell}(s_{t},\pi(\mathbf{o}_{t}))$ over $T$ timesteps, where $\gamma$ is a discount factor weighting the importance of future rewards.
MARL
We refer to a policy profile as $\pi=(\pi_{1},\ldots,\pi_{N})$, a policy profile excluding agent $i$ as $\pi_{-i}$, a joint policy profile for team $\mathcal{T}_{\ell}$ as $\pi_{\ell}=(\pi_{i},\forall i\in\mathcal{T}_{\ell})$, and a joint policy profile excluding all agents in $\mathcal{T}_{\ell}$ as $\pi_{-\ell}=(\pi_{i},\forall i\notin\mathcal{T}_{\ell})$. Each agent’s value function and state-action value (or $Q$) function characterise its expected returns under a policy profile $\pi$:
$V^{\pi_{i}}(s)=r_{i}+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,\pi_{1}(o_{1}),\ldots,\pi_{N}(o_{N}),s^{\prime})V^{\pi}_{i}(s^{\prime}),$
$Q^{\pi_{i}}(s,\mathbf{a})=r_{i}+\gamma\sum_{s^{\prime}\in S}P(s,a_{1},\ldots,a_{N},s^{\prime})V^{\pi_{i}}(s^{\prime})$
Our algorithm assumes access to value and $Q$-functions that take the global observations of all agents, not states, as input: $V^{\pi_{i}}(\mathbf{o})$ and $Q^{\pi_{i}}(\mathbf{o},\mathbf{a})$. Such $Q$-functions are used by actor-critic MARL algorithms that rely on centralised training with decentralised execution (Lowe et al. 2017; Foerster et al. 2018). Lastly, we define mean value functions and state-action value functions for each team: $\bar{V}^{\pi_{\ell}}(\mathbf{o}):=\frac{1}{|\mathcal{T}_{\ell}|}\sum_{i\in\mathcal{T}_{\ell}}V^{\pi_{i}}(\mathbf{o}),\bar{Q}^{\pi_{\ell}}(\mathbf{o},\mathbf{a}):=\frac{1}{|\mathcal{T}_{\ell}|}\sum_{i\in\mathcal{T}_{\ell}}Q^{\pi_{i}}(\mathbf{o},\mathbf{a})$.
Decision Trees
A decision tree (DT) recursively partitions an input space $\mathcal{X}$ through functions $f_{j}:\mathcal{X}\rightarrow\mathbb{R}$ and thresholds $\tau_{j}$ at each internal node $j$. Each internal node induces a partition of samples, $P_{j}={x\in\mathcal{X}:f_{j}(x)\leq\tau_{j}}$. For a DT policy, internal nodes $(f_{j},\tau_{j})$ encode observation-dependent decision criteria, while leaf nodes $l\in\mathcal{L}$ map partitioned observations to actions: $\hat{\pi}_{i}:\mathcal{O}_{i}\rightarrow\mathcal{A}_{i},\forall i\in[N]$.
VIPER
VIPER (Bastani, Pu, and Solar-Lezama 2018) is an imitation learning framework, adapted from the more general DAGGER (Ross, Gordon, and Bagnell 2011), that trains DTs as surrogate policies. Given a trained expert (NN) policy $\pi^{*}$, VIPER iteratively generates student (DT) policies $\hat{\pi}^{m}$. Specifically, in each iteration $m$, VIPER:
  1. (1)
    Collects $K$ new rollouts $\{\mathbf{o},\hat{\pi}^{m-1}(\mathbf{o})\}$ using the previous students from iteration $m-1$ (where $\hat{\pi}^{0}:=\pi^{*}$)
  2. (2)
    Resamples a dataset $\mathcal{D}$ from all trajectories collected so far, based on upweighting critical states where taking a suboptimal action may be costly in terms of $Q$-values:
    $p_{k}\propto V^{\pi^{*}}(\mathbf{o}_{k})-\min_{\mathbf{a}}Q^{\pi^{*}}(\mathbf{o}_{k},\mathbf{a})$
  3. (3)
    Relabels the dataset with the expert actions $\pi^{*}(\mathbf{o}_{k})$
  4. (4)
    Trains new DT students $\hat{\pi}^{m}$ on $\{\mathbf{o}_{k},\pi^{*}(\mathbf{o}_{k})\mid\mathbf{o}_{k}\in\mathcal{D}\}$
After $M$ iterations, VIPER (5) selects a student through validation on an additional set of rollouts. Ross, Gordon, and Bagnell (2011) showed that such a procedure is guaranteed to find a student that is performant on the distribution of states that it induces.
4 HYDRAVIPER
In this section, we present HYDRAVIPER (Algorithm 1), our algorithm for performant and efficient interpretable MARL. As shown in Figure 1, HYDRAVIPER builds on the DAGGER and VIPER frameworks by iteratively collecting data from environment rollouts to train DT policies. HYDRAVIPER first (1) partitions agents into clusters for scalability (line 4). Next, in each of $M$ iterations, HYDRAVIPER: (2) collects a dataset of rollouts from the environment, using an adaptive procedure (lines 6–7); (3) resamples the dataset to prioritise learning the correct actions in critical states, using team-based $Q$-values (lines 9–10); (4) and trains DTs based on these datasets (lines 11–12). After completing all $M$ training iterations, HYDRAVIPER (5) identifies the best-performing student for each agent, using a multi-armed bandit algorithm, and returns them as a policy profile (lines 13–14). We now describe each of these algorithm components in detail.
Algorithm 1 HYDRAVIPER
1:Input: Markov game $(\mathcal{S},\mathcal{A},P,R_{i},O_{i})$, experts $\pi^{*}$, expert $Q$-functions $Q^{\pi^{*}}$, per-iteration rollout count $K_{\mathrm{train}}$, rollout budgets $(B_{\mathrm{train}},B_{\mathrm{valid}})$, threshold $\epsilon$, iteration count $M$, scaling factor $c$, agent distance function $d$
2:Output: Trained students $\hat{\pi}$
3:Initialise dataset $\mathcal{D}\leftarrow\emptyset$, policies $\hat{\pi}_{i}^{0}\leftarrow\pi_{i}^{*},\forall i\in N$
4:Initialise rollout count $n_{\mathrm{train}}\leftarrow 0$
5:Initialise dropped rollout count $K_{\mathrm{drop}}\leftarrow\infty$
6$\triangleright$ Section 4.4: Agent Clustering
7:Cluster agents $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\leftarrow\textbf{Partition}(\Gamma,\pi^{*},d)$
8:for $m\in\{1,\dots,M\}$ do
9:     $\triangleright$ Section 4.2: Training Rollouts
10:     $\mathcal{D},n_{\mathrm{train}}\leftarrow$ TR-A $(\mathcal{D},\hat{\pi}^{m-1},m,K_{\mathrm{train}},B_{\mathrm{train}}$, $K_{\mathrm{drop}},n_{\mathrm{train}})$
11:     Reinitialise dropped rollout count $K_{\mathrm{drop}}\leftarrow\infty$
12:     for each team $\mathcal{T}_{\ell}\in\{\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\}$ do
13:         $\triangleright$ Section 4.1: Dataset Resampling
14:         $\mathcal{D}_{\ell}^{\prime},K_{\mathrm{drop}}^{\prime}\leftarrow$ C-Q$(\mathcal{D}_{\ell},\mathcal{T}_{\ell},\pi^{*},Q^{\pi^{*}},\epsilon)$
15:         $K_{\mathrm{drop}}\leftarrow\min(K_{\mathrm{drop}},K_{\mathrm{drop}}^{\prime})$
16:         for each agent $i\in\mathcal{T}_{\ell}$ do
17:              $\hat{\pi}_{i}^{m}\leftarrow\textbf{TrainDT}(\mathcal{D}_{\ell}^{\prime})$
18:for each team $\mathcal{T}_{\ell}\in\{1,\dots,L\}$ do
19:     $\triangleright$ Section 4.3: Validation Rollouts
20:     $\hat{\pi}_{i},\forall i\in\mathcal{T}_{\ell}\leftarrow$ VR-UCB($(\{\hat{\pi}^{m}_{\ell}\}_{m=1}^{M},\mathcal{T}_{\ell},B_{\mathrm{valid}},c)$
21:return $\hat{\pi}=(\hat{\pi}_{1},\dots,\hat{\pi}_{N})$
4.1 Dataset Resampling: Centralised-Q Weighting
Algorithm 2 Centralised-Q Resampling (C-Q)
1:Input: Dataset $\mathcal{D}_{\ell}$, teams $\mathcal{T}_{\ell}$, experts $\pi^{*}$, expert $Q$-functions $Q^{\pi^{*}}$, threshold $\epsilon$
2:Output: Resampled dataset $\mathcal{D}_{\ell}^{\prime}$, dropped rollout count $K_{\mathrm{drop}}$
3: Set weights for each $(\mathbf{o}_{k},\mathbf{a}_{k})\in\mathcal{D}_{\ell}$: $p_{\ell k}\leftarrow\bar{V}^{\pi_{\ell}^{*}}(\mathbf{o}_{k})-\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))$
4:Update
$K_{\mathrm{drop}}\leftarrow\min_{\ell}\left\lceil\frac{1}{T}|\{(\mathbf{o}_{k},\mathbf{a}_{k})\in\mathcal{D}_{\ell}\mid p_{\ell k}\leq\epsilon\}|\right\rceil$
5:Resample dataset: $\mathcal{D}_{\ell}^{\prime}\leftarrow\{(\mathbf{o}_{k},\mathbf{a}_{k})\sim p_{\ell k}\}$
6:return $\mathcal{D}_{\ell}^{\prime},K_{\mathrm{drop}}$
VIPER-based algorithms include a dataset resampling step (Algorithm 1, lines 9–10) so that students can focus their learning on more critical states. At a high level, they construct a training dataset by computing sampling weights over the aggregated dataset of environment rollouts, typically using some notion of value based on the expert $Q$-functions. Measuring value is straightforward in the single-agent setting, but — as we have mentioned — a key obstacle in multi-agent learning is efficient coordination among agents. To address this challenge, HYDRAVIPER induces coordination in the resampling step using a team-based notion of value (Algorithm 2), but trains DTs independently for each agent.
Specifically, HYDRAVIPER resamples the dataset for DT construction based on weights $p_{\ell k}$, which compute the relative importance of each sample for each team of agents $\mathcal{T}_{\ell}$ (Algorithm 2, line 1). Prior work computes this importance based on individual $Q$-functions, meaning that each agent must maintain its own dataset and induce coordination through (typically computationally expensive) joint training procedures. By contrast, we propose an intuitive change: HYDRAVIPER uses the mean of the expert $Q$-functions within each team of coordinated agents, $\bar{Q}^{\pi_{\ell}^{*}}:=\frac{1}{|\mathcal{T}_{\ell}|}\sum_{j\in\mathcal{T}_{\ell}}Q^{\pi_{j}^{*}}$, so as to prioritise samples according to their value to the team. Then, we compute the weights as the difference in value between the optimal joint team action and the worst-case joint team action. Intuitively, highly-weighted samples are those where coordinating on joint actions matters for performance. The weights are defined as:
$p_{\ell k}$ $\propto\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\pi^{*}(\mathbf{o}_{k}))-\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))$
$=\bar{V}^{\pi_{\ell}^{*}}(\mathbf{o}_{k})-\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k})).$ (1)
For further gains in sample efficiency, HYDRAVIPER does not compute $p_{\ell k}$ by enumerating joint actions over all agents present in the environment. Instead, it only enumerates possible joint actions $\mathbf{a}_{\ell}$ over the team and uses expert actions $\pi_{-\ell}^{*}(\mathbf{o}_{-\ell})$ for opponent agents. This novel resampling procedure eliminates the need for per-agent datasets in IVIPER and MAVIPER, allowing agents to prioritise the same critical states without computationally expensive joint training.
HYDRAVIPER uses the jointly sampled dataset to independently train DTs for each agent $i$ (Algorithm 1, lines 8–9). Each DT $\hat{\pi}_{i}$ uses individual observations $o_{i}$ to fit $\pi_{i}^{*}$’s actions in the dataset. Modifying the input dataset rather than the training procedure provides HYDRAVIPER with flexibility in the choice of DT learning algorithm. More advanced models such as random forests or mixtures of DTs (Vasić et al. 2022) can be incorporated to improve performance.
4.2 Training Rollouts: Adaptive Budget Allocation
Algorithm 3 Adaptive Training Rollouts (TR-A)
1:Input: Dataset $\mathcal{D}$, students $\hat{\pi}^{m-1}$, iteration $m$, per-iteration rollout count $K_{\mathrm{train}}$, training rollout budget $B_{\mathrm{train}}$, dropped rollout count $K_{\mathrm{drop}}$, total rollout count $n_{\mathrm{train}}$
2:Output: Updated dataset $\mathcal{D}$, total rollout count $n_{\mathrm{train}}$
3:Set $K_{\mathrm{train}}^{m}\leftarrow\min(K_{\mathrm{drop}},K_{\mathrm{train}})\mathds{1}[n_{\mathrm{train}}\leq B_{\mathrm{train}}]$
4:Update $n_{\mathrm{train}}\leftarrow n_{\mathrm{train}}+K_{\mathrm{train}}^{m}\mathds{1}[m>1]$
5:for each team $\mathcal{T}_{\ell}\in\{1,\dots,L\}$ do
6:      Collect and relabel $K_{\mathrm{train}}^{m}$ rollouts: $\mathcal{D}_{\ell}^{m}\leftarrow\{(\mathbf{o}_{\ell},\pi_{\ell}^{*}(\mathbf{o}_{\ell}))\sim d^{(\hat{\pi}^{m-1})}\}$
7:     Aggregate dataset: $\mathcal{D}_{\ell}\leftarrow\mathcal{D}_{\ell}\cup\mathcal{D}_{\ell}^{m}$
8:return $\mathcal{D},n_{\mathrm{train}}$
Thus far, we have assumed that HYDRAVIPER has access to a dataset of observation-action pairs for training. To collect this dataset, HYDRAVIPER follows the DAGGER-style iterative procedure of collecting a dataset at each iteration $m$ by rolling out the current student policies $\hat{\pi}^{m-1}$ (Algorithm 1, line 6–7). The next set of students are trained on the aggregate of all collected datasets, therefore building up the set of inputs likely to be encountered by the student policies during execution. However, collecting training rollouts is computationally expensive. Prior work employs an inefficient static allocation strategy which uniformly performs $K_{\textrm{train}}$ rollouts in each iteration. This strategy is problematic because the students are far from convergence early in training, so the distribution of trajectories collected earlier in training potentially diverges from those that converged students would encounter. HYDRAVIPER addresses this challenge through an adaptive rollout strategy that dynamically allocates the training budget at each iteration and prioritises critical states encountered later in training.
Recall that, for each team of cooperative agents $\mathcal{T}_{\ell}$, HYDRAVIPER follows Equation 1 to compute weights $p_{\ell k}$ for resampling the training dataset. We show the following:
Theorem 1.
Given a dataset of observation-action pairs for team $\mathcal{T}_{\ell}$ in iteration $m$ of HYDRAVIPER, $\mathcal{D}_{\ell}=\{(\mathbf{o}_{\ell},\mathbf{a}_{\ell})\}$, assume there exists a pair $(\mathbf{o}_{\ell k},\mathbf{a}_{\ell k})$ that receives the weight $p_{\ell k}^{(m)}=0$. Then, in iteration $m+1$ of HYDRAVIPER, this pair also receives the weight $p_{\ell k}^{(m+1)}=0$.
Proof.
See Appendix A.
As a result, samples $(\mathbf{o}_{\ell k},\mathbf{a}_{\ell k})$ with $p_{\ell k}=0$ are effectively removed from the dataset $\mathcal{D}$. This intuition serves as the motivation behind HYDRAVIPER’s adaptive training rollout budget allocation (Algorithm 3): after samples are dropped during the resampling procedure, HYDRAVIPER performs rollouts to replenish the dataset.
Specifically, we treat the first iteration as a warm-up period, in which the experts collect a predefined number of $K_{\mathrm{train}}$ rollouts (Algorithm 3, lines 3–5). This leads to an initial dataset of $T\cdot K_{\mathrm{train}}$ observation-action pairs. Each team $\mathcal{T}_{\ell}$ discards non-critical samples from its dataset (Algorithm 2, line 2), i.e. those where the range in the $Q$-value is at most a predefined threshold $\epsilon$. With the goal of efficiency in mind, HYDRAVIPER computes the minimum number of such discarded samples across all teams of cooperative agents. This then determines the minimum number of rollouts required to collect at least this many samples in the next iteration. As a result, the expected number of dropped rollouts, and therefore the budget for the next iteration, is:
$K_{\mathrm{drop}}=\min_{\ell}\left\lceil{\frac{1}{T}|\{(\mathbf{o}_{k},\mathbf{a}_{k})\in\mathcal{D}_{\ell}\mid p_{\ell k}\leq\epsilon\}|}\right\rceil.$
Over the remaining $M-1$ iterations, HYDRAVIPER continues collecting rollouts using students until it exhausts its total budget of $B_{\mathrm{train}}$ training rollouts (Algorithm 3, line 1). Choosing different rollout budgets allows performance and efficiency to be traded off. A higher budget is likely to lead to superior performance, as more rollouts will be collected from students closer to convergence before the budget is exhausted, but it also requires more computation time.
4.3 Validation Rollouts: UCB Policy Selection
Algorithm 4 UCB Validation Rollouts (VR-UCB)
1:Input: Policies $\{\hat{\pi}^{m}_{\ell}\}_{m=1}^{M}$, team $\mathcal{T}_{\ell}$, validation rollout budget $B_{\mathrm{valid}}$, scaling factor $c$
2:Output: Selected policies $\hat{\pi}_{i},\forall i\in\mathcal{T}_{\ell}$
3:Initialise $n_{m}\leftarrow 0$ for all $m\in\{1,\dots,M\}$
4:Initialise $n_{\min}\leftarrow\left\lceil{2\ln B_{\mathrm{valid}}}\right\rceil$
5: Initialise return estimates: $\mu_{\ell}^{m}\leftarrow\frac{1}{C_{min}}\sum_{k=1}^{C_{min}}\bar{R}_{\ell k},\bar{R}_{\ell k}\sim d^{(\hat{\pi}^{m}_{\ell},\pi^{*}_{-\ell})}$
6:for rollout $k\in\{1,\ldots,(B_{valid}-mn_{\min})\}$ do
7:     Set $m^{*}\leftarrow\operatorname{argmax}_{m}\hat{\mu}_{\ell}^{m}+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}}}$
8:     Collect mean return: $\bar{R}_{\ell k}\sim d^{(\hat{\pi}^{m^{*}}_{\ell},\pi^{*}_{-\ell})}$
9:     Update rollout count: $n_{m^{*}}\leftarrow n_{m^{*}}+1$
10:      Update running average of mean return: $\hat{\mu}_{\ell}^{m^{*}}\leftarrow\frac{n_{m^{*}}-1}{n_{m^{*}}}\hat{\mu}_{\ell}^{m^{*}}+\frac{1}{n_{m^{*}}}\bar{R}_{\ell k}$
11:return $\hat{\pi}_{\ell}^{m^{*}},\forall i\in\mathcal{T}_{\ell}$
Following $M$ iterations, HYDRAVIPER produces $M$ joint policy profiles for each team. It then needs to select the best-performing policy profile (Algorithm 1, lines 13–14). HYDRAVIPER iterates through the policy profiles to estimate the team performance of each one using a set of validation rollouts. The performance metric it uses is the undiscounted mean return of the team, $\bar{R}^{m}_{\ell}=\frac{1}{T}\sum_{t=0}^{T}R_{\ell}(s_{t},\hat{\pi}^{m}_{\ell}(\mathbf{o}_{t}))$.
As is the case for training, collecting validation rollouts is computationally intensive, so these rollouts also need to be efficiently allocated. However, the problem setting differs here. Our goal is not to collect a diverse set of training rollouts, but rather to identify the most performant policy profiles using as few rollouts as possible. The mean return of each policy profile is unknown a priori; it must be estimated by selecting policy profiles and performing rollouts with noisy returns. Again, a fixed allocation strategy of $K_{\mathrm{valid}}$ environment rollouts for each policy profile is wasteful. The rollouts assigned to clearly badly-performing policy profiles could be reallocated to reduce the variance in the estimated returns of promising policy profiles. This motivation aligns with that of multi-armed bandit (MAB) problems.
Given a limited budget of $B_{\mathrm{valid}}$ rollouts, we represent the task of selecting the best-performing policy profile as a MAB problem. For each team $\mathcal{T}_{\ell}$, the policy profile $\hat{\pi}_{\ell}^{m}$ from each iteration $m$ is an arm, and its return is a random variable $\bar{R}_{\ell}^{m}$ with unknown mean $\mu_{\ell}^{m}$. Each rollout samples from one of these random variables, which captures the distribution of returns from environment and policy randomness. The objective is to identify the best arm $m_{\ell}^{*}=\operatorname{argmax}_{m}\mu_{\ell}^{m}$ in as few rollouts as possible, i.e. to minimise the regret with respect to the policy that selects $m_{\ell}^{*}$ for every rollout.
In this work, we use a modification of the UCB1 algorithm (Auer, Cesa-Bianchi, and Fischer 2002). This allows us to achieve logarithmic regret given a readily satisfiable assumption: that the returns $\bar{R}_{\ell}^{m}$ of the arms are bounded. Although UCB assumes that the arms are bounded in $[0,1]$, it can be modified in a manner equivalent to rescaling the rewards to remain in $[0,1]$. We rely on the general form of the Chernoff-Hoeffding bound:
Theorem 2.
(Theorem 2 of Hoeffding (1963)) For independent random variables $X_{1},\ldots,X_{n}$ with mean $\mu$, $\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$, and $a_{i}\leq X_{i}\leq b_{i},\forall i\in\{1,\ldots,n\}$, then for $\alpha>0$
$\Pr(\bar{X}\geq\mu+\alpha)\leq e^{-\frac{2n^{2}\alpha^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}}}.$
Corollary 1.
Assume that $\bar{R}_{\ell}^{m}$ is bounded by $[a,b]$ with $\Delta=b-a$ for all $i,m$. For $c=2\Delta^{2}$, Theorem 2 shows that, for the empirical mean $\hat{\mu}_{\ell}^{m}$, (notation simplified for clarity)
$\Pr\left(\hat{\mu}_{\ell}^{m}\geq\mu_{\ell}^{m}+\sqrt{\frac{c\ln B}{n_{m}}}\right)\leq e^{-\frac{2c\ln B}{\Delta^{2}}}=B^{-4}.$
This is the same bound demonstrated for UCB1, and the same holds for the lower confidence bound $\mu_{\ell}^{m}-\sqrt{\frac{c\ln B}{n_{m}}}$. Overall, this choice of $c$ leads to the same $O(\log B)$ regret bound as UCB1. However, HYDRAVIPER can also be extended to use other MAB algorithms. If the mean returns of each policy profile are assumed to be normally distributed, the UCB1-Normal algorithm (Auer, Cesa-Bianchi, and Fischer 2002) could be used; it also achieves logarithmic regret. This algorithm effectively chooses $c$ to be proportional to the arms’ sample variance; the greater the variance, the wider the confidence bound. It is also possible to obtain an offline (but biased) estimate of the sample variance by performing expert rollouts before running UCB.
Given a total budget of $B_{\mathrm{valid}}$ validation rollouts, HYDRAVIPER performs them as follows. For each policy profile, it first performs $n_{\min}=\left\lceil{2\ln B_{\mathrm{valid}}}\right\rceil$ rollouts to generate initial estimates of the mean returns (Algorithm 4, lines 2–3). To allocate the remainder of the budget (lines 4–8), HYDRAVIPER follows UCB1 to select the policy profile index for the $k$th validation rollout as
$m_{\ell k}^{*}=\operatorname{argmax}_{m}\left(\hat{\mu}_{\ell}^{m}(k)+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}(k)}}\right),$
where $n_{m}(k)=\sum_{k^{\prime}=1}^{k}\mathds{1}[m_{\ell k^{\prime}}^{*}=m]$ is the number of rollouts that have used policy profile $m$ thus far, $\hat{\mu}_{\ell}^{m}(k)=\frac{\sum_{k^{\prime}=1}^{k}\bar{R}_{\ell k^{\prime}}\mathds{1}[m_{\ell k^{\prime}}^{*}=m]}{n_{m}(k)}$ is the empirical mean of the returns $\bar{R}_{\ell k}$ from policy profile $m$, $B_{\mathrm{valid}}$ is the total budget of rollouts, and $c$ is a scaling constant for the confidence bound (see Section 5.4). HYDRAVIPER maintains a running average for the mean return of each policy profile, which it updates using the mean return $\bar{R}_{\ell k}$ of each rollout (line 8).
4.4 Agent Clustering: Scaling Up HYDRAVIPER
When resampling the dataset, HYDRAVIPER calculates sample weights following Equation 1. This computation requires enumerating joint actions $\mathbf{a}_{\ell}$ for each team $\mathcal{T}_{\ell}$, in order to find the worst-case joint action that minimises the team-averaged $Q$-function, $\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))$. The complexity of this step scales with the size of the joint action space, and thus exponentially with the size of the team. Some mixed competitive-cooperative environments (see Section 5.1) have inherent team structure that can reduce this complexity. In cooperative environments such as TSC, HYDRAVIPER clusters the agent set into teams to improve training efficiency (Algorithm 3).
Algorithm 5 Agent Graph Clustering (Partition)
1:Input: Markov game $(\mathcal{S},\mathcal{A},P,R_{i},O_{i})$, experts $\pi^{*}$, agent distance function $d$
2:Output: Agent teams $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}$
3:Construct graph $G=(V=\{1,\ldots,N\},E,w=0)$
4:for each agent $i\in\{1,\dots,N\}$ do
5:     for each agent $j\in\{1,\dots,N\}$ do
6:         Assign edge weight $w_{ij}\leftarrow\frac{1}{d(i,j)}$
7:Partition graph $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\leftarrow\textbf{METIS}(G,L)$
8:return $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}$
Our goal is to find a clustering of the agent set into teams $\mathcal{T}_{1}\ldots\mathcal{T}_{L}$, so that HYDRAVIPER-trained DT students have performance similar to training on the full agent set, but improved scalability in that the number of actions to enumerate per team is much smaller than the full agent set: $\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\ll\prod_{i\in\{1,\ldots,N\}}\mathcal{A}_{i}$. We leverage the intuition that agents distant from each other (in terms of environmental distance, trajectory similarity, or other metrics) are unlikely to be influential on each other in most environments.
Suppose that we are given a function $d(i,j)$ that computes this distance between a pair of agents. In our clustering procedure (Algorithm 5), we first construct a complete graph $G=(V,E)=K_{N}$ where the nodes represent agents, and the weight between node $i$ and node $j$, $w_{ij}$, is inversely proportional to $d(i,j)$ (lines 1–4). Then, we perform graph partitioning to divide $G$ into $L$ contiguous, connected node clusters of approximately equal size (line 5), such that the sum of the weights of inter-cluster edges is minimised. We use the hierarchical METIS algorithm (Karypis and Kumar 1998) to accomplish this. Note that we solve a graph partitioning problem instead of a min-cut problem to prevent the clusters from being imbalanced. Otherwise, in the worst case, the largest cluster could have size $O(N)$, thus yielding minimal gains in scalability.
How can the distance metric $d$ be defined? For an environment that has inherent team structure (such as the physical deception environment in Section 5.1), we define the graph $G$ to consist of a complete subgraph for each team:
$d_{ij}=\begin{cases}1&\textrm{if $i$ and $j$ share a team}\\ \infty&\textrm{otherwise}\end{cases}$
For traffic signal control environments, we note that the road network inherently forms a graph $G_{env}$, which can be partitioned to obtain sets of spatially proximal agents that correspond to neighbouring intersections. In this case, we simply set $G=G_{env}$, or equivalently
$d_{ij}=\begin{cases}1&\textrm{$e_{ij}\in G_{env}$}\\ \infty&\textrm{$e_{ij}\notin G_{env}$}\end{cases}$
5 Experiments
Now, we demonstrate the utility of HYDRAVIPER for interpretable MARL using experiments on various benchmark environments. In doing so, we perform a functionally grounded evaluation of interpretability (Doshi-Velez and Kim 2017), where we assess the quality of the generated DTs in terms of performance and computational efficiency. As the DTs would be used directly in lieu of NN-based policies in deployment, we consider these to be good proxy metrics of their practical applicability. More specifically, we address the following research questions:
RQ 1.
Is HYDRAVIPER both performant and efficient (in terms of environment interactions and runtime)?
RQ 2.
Does HYDRAVIPER maintain performance optimality as the environment interaction budget decreases?
RQ 3.
Can HYDRAVIPER maintain performance optimality while scalability is improved through agent clustering?
5.1 Environments
We evaluate HYDRAVIPER on four environments: two environments in the multi-agent particle world (MPE) benchmark (Lowe et al. 2017), and two traffic signal control (TSC) environments in the RESCO benchmark (Ault and Sharon 2021). In MPE environments, agents must navigate in a 2D space to accomplish a coordinated objective, making these environments ideal for assessing coordination capabilities.
Cooperative navigation (CN)
In this environment, a team of three agents must coordinate to split up and cover three different targets while avoiding collisions with each other.
Physical deception (PD)
In this environment, a team of two defender agents must cooperate to protect two targets from an adversary agent. One of the two targets is the “goal” for the adversary; this is not known to the adversary, which can only observe the positions of the targets and defenders. We train the two defender agents against an NN adversary.
In TSC environments, each agent controls a single intersection by selecting different signal phases; each phase permits vehicles from a subset of lanes to pass through the intersection. Both environments are based on real-world road corridors reproduced in the traffic simulator SUMO (Alvarez Lopez et al. 2018). To interface with the simulator, we use the OpenAI Gym-style wrapper sumo-rl (Alegre, Bazzan, and da Silva 2021). We focus on imitating experts for all agents as a team.
Cologne corridor (CC)
(Uppoor and Fiore 2011) This environment simulates three signalised intersections in a corridor from the city of Cologne (Köln), Germany. It has a total volume of 4 494 vehicles in 7–8 am rush hour traffic.
Ingolstadt corridor (IC)
(Lobo et al. 2020) This larger environment simulates seven signalised intersections in a corridor from the city of Ingolstadt, Germany. It has a lower total volume of 3 031 vehicles in 4–5 pm rush hour traffic.
5.2 Baselines and Setup
We compare HYDRAVIPER with IVIPER and MAVIPER, which represent the state of the art in interpretable multi-agent RL with DT surrogate policies. We additionally compare with expert policies — MADDPG (Lowe et al. 2017) for MPE and MPLight (Chen et al. 2020) for TSC — and an additional baseline, imitation DT. Imitation DT does not use students to collect rollouts, nor does it perform dataset resampling; it collects the same number of training rollouts as the other algorithms and trains DTs on the collected dataset. As imitation DT performs worse than the other algorithms by a wide margin, we do not include it in Table 1 or Figure 2 but show its performance in Appendix B.
For MPE environments, we use a horizon of 25 timesteps per episode, and we trained MADDPG for 60 000 episodes as the expert for the DT students to imitate. For TSC environments, we use a horizon of 125 timesteps per episode (each timestep represents 20 seconds of simulation time), and we trained MPLight for 500 episodes as the expert. All imitation learning algorithms were run for 100 iterations to produce DTs with a maximum depth of 4. IVIPER and MAVIPER ran $K_{\mathrm{train}}=K_{\mathrm{valid}}=50$ training and validation rollouts per iteration for MPE (including for the initial iteration where rollouts are collected by the experts), and 10 rollouts per iteration for TSC. Imitation DT ran the same numbers of training rollouts. We set these to equalise the number of environment interactions per iteration.
We repeated all experiments 10 times with different random seeds, and we report the mean and 95% confidence interval of the reward over 10 rollouts performed with the final student policy profiles generated from these runs. Most experiments were run in parallel on a server with 56 2.75GHz AMD EPYC 7453 processors and 252 GiB of RAM. For these experiments, we report the number of rollouts collected, not runtimes, as the time per rollout is roughly constant. However, we also report runtimes for the execution of IVIPER, MAVIPER, and HYDRAVIPER on all four environments. For these experiments, we use the kernprof profiler (v4.1.3) to run them in sequence, with no other concurrent processes running except system routines. These experiments were run on another server with 8 4.2GHz Intel i7-7700K processors and 62 GiB of RAM.
5.3 Results
Environment Expert IVIPER MAVIPER HYDRAVIPER HYDRAVIPER (LB)
Cooperative Navigation Total Penalty 122.67 $\pm$ 1.67 160.87 $\pm$ 4.31 144.35 $\pm$ 2.12 144.48 $\pm$ 2.67 144.84 $\pm$ 2.12
Runtime (s) N/A 2 444.6 $\pm$ 9.1 21 188.7 $\pm$ 408.6 206.2 $\pm$ 11.1 180.5 $\pm$ 9.3
Physical
Deception
Defender
Reward
8.19 $\pm$ 0.50 6.94 $\pm$ 0.52 7.74 $\pm$ 0.82 7.72 $\pm$ 0.53 7.12 $\pm$ 0.84
Runtime (s) N/A 2 017.2 $\pm$ 21.3 11 782.4 $\pm$ 137.8 1 173.5 $\pm$ 21.6 388.4 $\pm$ 5.6
Cologne Corridor Queue Length 18.94 $\pm$ 2.49 22.06 $\pm$ 2.91 25.85 $\pm$ 5.22 16.72 $\pm$ 1.51 18.77 $\pm$ 3.69
Runtime (s) N/A 33 841.4 $\pm$ 441.3 37 503.8 $\pm$ 834.8 13 651.6 $\pm$ 254.2 1 865.4 $\pm$ 26.1
Ingolstadt Corridor Queue Length 23.01 $\pm$ 1.10 21.51 $\pm$ 2.13 24.26 $\pm$ 2.54 19.77 $\pm$ 1.51 21.87 $\pm$ 1.59
Runtime (s) N/A 75 709.6 $\pm$ 441.3 52 316.9 $\pm$ 6 280.2 11 203.9 $\pm$ 67.2 6 490.0 $\pm$ 98.0
Table 1: Performance and runtimes (means and 95% confidence intervals) for HYDRAVIPER and baselines. All algorithms are given the same environment interaction budget, except for low-budget (LB) HYDRAVIPER (which uses 20% of the rollouts for MPE, 10% of the rollouts for TSC). HYDRAVIPER achieves or exceeds the performance of MAVIPER using a fraction of the runtime, and still performs well in the low-budget setting. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Appendix C shows runtimes for individual algorithm steps.
RQ 1
HYDRAVIPER achieves strong, coordinated performance in a computationally efficient manner. First, we assess HYDRAVIPER’s performance as we vary it between two environment interaction budget levels, high (5 000 training/5 000 validation rollouts for MPE, 1 000 training/1 000 validation rollouts for TSC) and low (500 training/1 500 validation rollouts for MPE, 100 training/100 validation rollouts for TSC). As shown in Table 1, HYDRAVIPER students perform better than or comparable to students trained by the most performant DT baseline (MAVIPER for MPE, IVIPER for TSC) on all environments at both budget levels. HYDRAVIPER’s performance is also better than or comparable to the NN experts for all environments except cooperative navigation, in which all DT-based algorithms cannot achieve expert-level performance. In physical deception, although neither MAVIPER nor HYDRAVIPER substantially outperforms IVIPER given the considerable stochasticity in the environment, HYDRAVIPER achieves a level of performance much closer to MAVIPER, while its training time is an order of magnitude less than MAVIPER.
In the TSC environments, HYDRAVIPER is the best-performing algorithm at both high and low interaction budget levels. Notably, HYDRAVIPER at the high budget level substantially outperforms the expert on the Ingolstadt corridor. By contrast, MAVIPER fails to coordinate the intersection agents and is overall the worst-performing algorithm. HYDRAVIPER more than halves the runtime from both IVIPER and MAVIPER on both TSC environments.
Figure 2: Performance of HYDRAVIPER and baselines as the number of rollouts decreases. Top shows decreasing training rollouts; bottom shows decreasing validation rollouts. HYDRAVIPER’s performance stays consistent as the number of rollouts decreases. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Bars show 95% confidence intervals based on 10 randomly-seeded runs. Full results are shown in Table 2 in Appendix B.
RQ 2
As the environment interaction budget decreases, HYDRAVIPER still outperforms baselines. Now, we investigate the ability of HYDRAVIPER to adapt to increasing budget constraints for environment interaction, as would be imposed by users who wish to quickly iterate on DT policy training. As shown in Figure 2, HYDRAVIPER’s performance in all four environments does not substantially change as the training and validation rollout budgets are reduced individually. Furthermore, in all four environments, HYDRAVIPER achieves performance on par with or better than MAVIPER at all budget levels. Therefore, HYDRAVIPER is able to maintain a Pareto frontier in the tradeoff between performance and computational efficiency.
In cooperative navigation, the performance of both HYDRAVIPER and MAVIPER remain similar as the training and validation budgets are reduced individually. However, when both budgets are reduced simultaneously (shown in Figure 5), the performance of HYDRAVIPER but not MAVIPER remains essentially unchanged. In physical deception, HYDRAVIPER still performs well even as its training budget is reduced by a factor of 10, whereas MAVIPER performs substantially worse. Furthermore, the 95% confidence intervals of HYDRAVIPER’s rewards are smaller than MAVIPER at all validation budget levels. Thus, HYDRAVIPER is able to identify performant policy profiles more consistently than MAVIPER.
In the Cologne corridor, HYDRAVIPER’s performance consistently remains within the expert’s 95% confidence interval at all environment interaction budget levels, whereas the same is not true of MAVIPER. Meanwhile, the performance of HYDRAVIPER on the Ingolstadt corridor substantially exceeds the expert at all budget levels, whereas MAVIPER and IVIPER (except for the 500 validation rollout setting) remain in the expert’s 95% confidence interval.
RQ 3
Even when the agent set is decomposed through clustering, HYDRAVIPER maintains its performance. Finally, we evaluate the effect of agent clustering on the performance of HYDRAVIPER. For physical deception, HYDRAVIPER by default runs with the adversary and the two defenders on opposing teams, and Table 1 and Figure 2 show that it achieves good performance. For the Ingolstadt corridor environment in the low-budget setting (100 training/100 validation rollouts), we evaluate the effect of directly clustering the agent set into two teams based on the road network graph $G_{env}$. As shown in Figure 3, the graph partitioning method introduced in Section 4.4 allows HYDRAVIPER to retain its performance even when the size of the agent set is approximately halved for each team. This reduces the size of the joint action set for the largest team from 1 944 to 72, and the total runtime of HYDRAVIPER by 20%. This is a notable improvement given the relatively small size of the dataset in the low-budget setting.
Our graph partitioning method also outperforms two baselines. First, its mean reward and reward variance are both much smaller than a random agent clustering method, in which agents are randomly assigned to one of two teams. Second, graph partitioning delivers a small performance improvement over the contiguous clustering method, which randomises between all possible cuts of the road network graph into two connected components. The strong performance of this baseline suggests that incorporating environmental structure is important to a high-quality clustering. Unlike the two baselines, however, our graph partitioning method always yields a clustering that is evenly distributed in size. The resulting clustering also corresponds to the cut under the contiguous method with the best performance.
Clustering Method Runtime (s) Worst-Performing
Clustering
HYDRAVIPER LB 6 490.0 $\pm$ 98.0 s [[1,2,3,4,5,6,7]]
+ random 5 130.7 $\pm$ 104.2 s [[1,2,6],[3,4,5,7]]
+ contiguous 5 135.6 $\pm$ 99.0 s [[1,2,3,4,5],[6,7]]
+ graph-partition 5 181.8 $\pm$ 149.3 s [[1,2,3],[4,5,6,7]]
Figure 3: (L) Performance of HYDRAVIPER on the Ingolstadt corridor (IC) under different agent clustering methods at the low budget level. (R) Runtimes and resulting clusterings (worst-performing clustering across 10 different random seeds shown) of HYDRAVIPER under these methods. For the clusterings, the intersection agent numbers follow those shown in Figure 1. Note that graph-partition achieves the best performance and runtime.
5.4 Hyperparameter Sensitivity
Figure 4: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the cooperative navigation (CN) environment. Lower rewards are better. The default value is 4 for both hyperparameters.
To understand the effects of HYDRAVIPER’s hyperparameters on its performance, we conduct experiments to vary the depth of the DT students, and the scaling constant $c$ for UCB policy selection (Section 4.3), on the cooperative navigation environment. We choose this environment due to its relatively low level of randomness. For these experiments, we use HYDRAVIPER at the low budget level (500 training/1 500 validation rollouts) as the baseline algorithm, and fix all hyperparameters other than those of interest.
By default, we use a DT depth of 4; our results show that DTs of this depth provide a good tradeoff between expressiveness and computational efficiency. As shown on the left of Figure 4, depth-4 DTs outperform depth-2 and depth-3 DTs on cooperative navigation. This is an intuitive result; the optimal agent policy for this environment cannot be represented with such shallow DTs, as they must condition on the positions of the other agents and the landmarks. However, we find that depth-4 DTs also marginally outperform depth-5 DTs. This same pattern exists in all of the environments that we use for evaluation. We hypothesise that the amount of data collected by HYDRAVIPER at the low budget level is insufficient to coordinate between depth-5 DTs.
Our default value for $c$ is also 4. As we outlined in Corollary 1 in Section 4.3, the value that theoretically allows logarithmic regret to be achieved is based on the range of mean returns in the environment. Since the agents navigate in a $2\times 2$ square environment, the maximum distance of an agent to a target is $2\sqrt{2}$. The reward in this environment is the negation of the minimum agent distance to each landmark, plus a penalty of -1 for each agent that the ego agent collides with. Therefore, the maximum possible penalty is $\Delta=3\cdot 2\sqrt{2}+2$, which requires $c\approx 219.88$ to achieve the guarantee of Corollary 1. However, as shown on the right of Figure 4, $c=4$ empirically performs best; $c=8$ and $c=16$ are already excessively conservative given the low randomness in the environment. Since the agents are already trained, we hypothesise that the potential range of returns is less useful in practice for finding a good policy profile than the typical range of returns. This can be approximated by performing expert rollouts, as we described in Section 4.3.
Sensitivity results for the three other environments are shown in Appendix D. The best DT depths for these environments are all 4, as with cooperative navigation, while the best tested values of $c$ for physical deception and the Cologne corridor are respectively 2 and 16.
Figure 5: Ablation of HYDRAVIPER on the cooperative navigation (CN, left) and Cologne corridor (CC, right) environments. Lower rewards are better.
5.5 Ablation
Lastly, to understand which components of HYDRAVIPER are responsible for its success, we conduct an ablation study for HYDRAVIPER at two budget levels in the cooperative navigation (CN) and Cologne corridor (CC) environments. For the high budget level, we use 5 000 training/5 000 validation rollouts for CN, and 1 000 training/1 000 validation rollouts for CC; for the low budget level, we use 500 training/1 500 validation rollouts for CN, and 100 training/100 validation rollouts for CC. We compare HYDRAVIPER’s centralised-$Q$ resampling with IVIPER’s independent resampling (HYDRAVIPER - CQ). Also, we study the impact of removing adaptive training budget allocation (HYDRAVIPER - TR-A) and UCB-based validation budget allocation (HYDRAVIPER - VR-UCB). Results for physical deception and the Ingolstadt corridor are shown in Appendix E.
Figure 5 shows the results of our ablation study. On both environments and at both budget levels, centralised-$Q$ resampling outperforms the IVIPER resampling scheme, although the gap in performance is less pronounced for the Cologne corridor due to environmental randomness. This result suggests that sampling the training dataset independently for each agent, instead of according to team performance, is insufficient for achieving coordinated behaviour in the resulting students. Meanwhile, removing the budget allocation methods degrades the performance of HYDRAVIPER. Having either one of the budget allocation methods is generally sufficient to improve HYDRAVIPER’s reward, except in one case: for cooperative navigation at the low budget level, HYDRAVIPER performs worse when only one budget allocation mechanism is present. Meanwhile, for the Cologne corridor at the low budget level, the variance in HYDRAVIPER’s reward is large both when only centralised-$Q$ resampling is present, and also when it is removed. These results suggest that the primary benefit of the two rollout budget allocation mechanisms is to stabilise HYDRAVIPER’s learning process, especially in the low budget setting when extracting the most information from each rollout is critical.
6 Conclusion and Future Work
In this work, we introduced a new DT-based interpretable MARL method, HYDRAVIPER. HYDRAVIPER addresses several limitations of prior multi-agent methods that follow the VIPER framework: (1) it improves performance by using a joint dataset resampling scheme based on team $Q$-values, and (2) it improves computational efficiency by adaptively allocating fixed budgets of environment interactions for training and validation, as well as by dividing agents into jointly-trained teams. Based on experiments in benchmark environments for multi-agent coordination and traffic signal control, we showed that HYDRAVIPER achieves performance comparable with MAVIPER (a centralised method) and even neural network experts, all within a runtime less than IVIPER (a decentralised method). We also demonstrated HYDRAVIPER’s sample efficiency in its ability to retain a similar level of performance using a fraction of the environment interactions.
Through our experiments in the Ingolstadt corridor environment, we scaled up the VIPER framework to seven agents. To our knowledge, this is the largest team of coordinated agents to which interpretable MARL has been applied thus far. However, environments based on real-world domains can have many more agents than the environments that we studied. For instance, the review of Noaeen et al. (2022) showed that TSC environments of dozens or even hundreds of agents are used in the RL literature. At the extreme, Chen et al. (2020) used parameter-shared MPLight policies as controller agents for an extremely large simulation of 2 510 traffic lights. Our agent clustering approach shows promise in scaling up to larger environments while retaining performance comparable to that of expert policies. In scaling up, other methods of assigning edge weights $e_{ij}$ to agent pairs could be developed to leverage structure present in MARL environments, including weights based on $Q$-values or on observation-action trajectories. We envision that the flexibility of the HYDRAVIPER framework will allow it to adapt to characteristics of different MARL environments while maintaining Pareto optimality in the performance-computational efficiency tradeoff.
Appendix A Proof of Resampling Theorem
Here, we provide a proof for Theorem 1 in Section 4.2.
Theorem 1.
Given a dataset of observation-action pairs for team $\mathcal{T}_{\ell}$ in iteration $m$ of HYDRAVIPER, $\mathcal{D}_{\ell}=\{(\mathbf{o}_{\ell},\mathbf{a}_{\ell})\}$, assume there exists a pair $(\mathbf{o}_{\ell k},\mathbf{a}_{\ell k})$ that receives the weight $p_{\ell k}^{(m)}=0$. Then, in iteration $m+1$ of HYDRAVIPER, this pair also receives the weight $p_{\ell k}^{(m+1)}=0$.
Proof.
If $p_{\ell k}^{(m)}\propto\bar{V}^{\pi_{\ell}^{*}}(\mathbf{o}_{k})-\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))=0$, then by definition
$\begin{split}\bar{V}^{\pi_{\ell}^{*}}(\mathbf{o}_{k}):=\max_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))=\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k})),\end{split}$ (2)
i.e. joint team actions $\mathbf{a}_{\ell}$ have no effect on the value given observation $\mathbf{o}_{k}$. When HYDRAVIPER resamples the dataset in iteration $m$ (Algorithm 2, line 3), $(\mathbf{o}_{\ell k},\mathbf{a}_{\ell k})$ will not be part of the resampled dataset $\mathcal{D}_{\ell}^{\prime}$. However, the resampled dataset $\mathcal{D}_{\ell}^{\prime}$ does not replace the original dataset $\mathcal{D}_{\ell}$.
In iteration $m+1$, $\mathcal{D}_{\ell}$ is aggregated with a newly-collected dataset of observation-action pairs $\mathcal{D}_{\ell}^{m}$ (Algorithm 3, line 5), and $(\mathbf{o}_{\ell k},\mathbf{a}_{\ell k})$ continues to be part of this dataset. A new set of weights $p_{\ell}^{(m+1)}$ are computed using this expanded dataset (Algorithm 1, line 8). Assume that $p_{\ell k}^{(m+1)}\neq 0$. Without loss of generality, let $p_{\ell k}^{(m+1)}>0$. Then by definition
$\bar{V}^{\pi_{\ell}^{*}}(\mathbf{o}_{k}):=\max_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))>\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k})).$
None of $\mathbf{o}_{k}$, $\pi^{*}_{-\ell}$, $\bar{V}^{\pi_{\ell}^{*}}$, or $\bar{Q}^{\pi_{\ell}^{*}}$ changed between iteration $m$ and $m+1$, since HYDRAVIPER uses experts and expert value functions to compute the weights. This contradicts Equation 2. Thus, if $p_{\ell k}^{(m)}=0$, then $p_{\ell k}^{(m+1)}=0$. ∎
A similar proof holds if the strict equality is replaced by the threshold $\epsilon$.
Appendix B Full Performance Results
In Table 2, we show the mean and 95% confidence intervals of the performance of IVIPER, MAVIPER, HYDRAVIPER, the expert, and imitation DT, on all four environments with all of the training and validation rollout budgets we tested.
Appendix C Full Runtime Results
In Figure 6 and Figure 7, we show the mean and 95% confidence intervals of the runtimes of IVIPER, MAVIPER, and HYDRAVIPER on the four most costly algorithm steps: training rollouts, dataset resampling, DT training, and validation rollouts. Figure 6 shows the runtimes for the multi-agent particle world environments, while Figure 7 shows the runtimes for the traffic signal control environments.
Environment Training Validation Expert Imitation DT IVIPER MAVIPER HYDRAVIPER
Cooperative 5000 5000 122.67 $\pm$ 1.67 221.19 $\pm$ 8.58 160.87 $\pm$ 4.31 144.35 $\pm$ 2.12 144.48 $\pm$ 2.67
Navigation 2500 5000 211.84 $\pm$ 6.32 161.62 $\pm$ 3.66 146.28 $\pm$ 3.34 144.13 $\pm$ 1.59
500 5000 218.22 $\pm$ 5.90 166.11 $\pm$ 2.67 146.91 $\pm$ 3.90 148.71 $\pm$ 2.80
5000 2500 161.94 $\pm$ 4.93 145.27 $\pm$ 3.07 144.66 $\pm$ 1.62
5000 1500 163.25 $\pm$ 3.60 145.10 $\pm$ 4.69 143.86 $\pm$ 1.54
500 1500 168.31 $\pm$ 5.52 153.66 $\pm$ 4.07 144.84 $\pm$ 2.12
Physical 5000 5000 8.19 $\pm$ 0.50 6.27 $\pm$ 0.43 6.94 $\pm$ 0.52 7.74 $\pm$ 0.82 7.72 $\pm$ 0.53
Deception 2500 5000 5.73 $\pm$ 0.31 6.60 $\pm$ 0.67 7.84 $\pm$ 0.66 7.91 $\pm$ 0.40
500 5000 5.32 $\pm$ 0.61 6.40 $\pm$ 0.44 6.30 $\pm$ 0.45 7.58 $\pm$ 0.91
5000 2500 7.51 $\pm$ 0.81 6.84 $\pm$ 1.21 7.38 $\pm$ 0.63
5000 1500 6.28 $\pm$ 0.80 6.64 $\pm$ 1.37 6.85 $\pm$ 0.70
500 1500 6.03 $\pm$ 0.70 7.36 $\pm$ 0.99 7.12 $\pm$ 0.84
Cologne 1000 1000 18.94 $\pm$ 2.49 137.67 $\pm$ 0.64 22.06 $\pm$ 2.91 25.85 $\pm$ 5.22 16.72 $\pm$ 1.51
Corridor 500 1000 211.84 $\pm$ 8.00 21.60 $\pm$ 2.50 28.82 $\pm$ 6.86 19.13 $\pm$ 2.70
100 1000 218.22 $\pm$ 10.29 22.05 $\pm$ 2.72 24.47 $\pm$ 5.21 16.75 $\pm$ 1.85
1000 500 22.07 $\pm$ 3.83 21.84 $\pm$ 4.33 20.06 $\pm$ 4.15
1000 100 19.73 $\pm$ 1.43 22.50 $\pm$ 4.38 17.12 $\pm$ 2.07
100 100 23.40 $\pm$ 3.53 28.91 $\pm$ 7.27 18.77 $\pm$ 3.69
Ingolstadt 1000 1000 23.01 $\pm$ 1.10 169.43 $\pm$ 5.26 21.51 $\pm$ 2.91 24.26 $\pm$ 2.54 19.77 $\pm$ 1.51
Corridor 500 1000 170.55 $\pm$ 3.41 21.79 $\pm$ 1.67 23.21 $\pm$ 1.12 18.48 $\pm$ 0.74
100 1000 166.99 $\pm$ 3.67 22.46 $\pm$ 2.52 25.20 $\pm$ 2.29 20.11 $\pm$ 1.57
1000 500 20.08 $\pm$ 0.97 24.30 $\pm$ 3.79 20.25 $\pm$ 1.26
1000 100 22.26 $\pm$ 1.62 22.18 $\pm$ 0.87 20.28 $\pm$ 2.04
100 100 27.23 $\pm$ 5.31 23.31 $\pm$ 1.44 21.87 $\pm$ 1.59
Table 2: Performance (means and 95% confidence intervals) for IVIPER, MAVIPER, and HYDRAVIPER at different training and validation budget levels. For physical deception, higher rewards are better; for other environments, lower rewards are better.
Figure 6: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the two multi-agent particle world environments, physical deception and cooperative navigation. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Figure 7: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the two traffic signal control environments, Cologne corridor and Ingolstadt corridor. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Appendix D Full Hyperparameter Sensitivity Results
In Figure 8, Figure 9, and Figure 10, we show the performance of HYDRAVIPER in the low-budget setting as two hyperparameters — DT depth and the UCB scaling constant $c$ — are varied for the physical deception, cooperative navigation, and Ingolstadt corridor environments.
Figure 8: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the physical deception (PD) environment. Higher rewards are better.
Figure 9: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the cooperative navigation (CN) environment. Lower rewards are better. The default value is 4 for both hyperparameters.
Figure 10: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the Ingolstadt corridor (IC) environment. Lower rewards are better.
Appendix E Full Ablation Results
In Figure 11, we show the performance of HYDRAVIPER on the physical deception (left) and Ingolstadt corridor (right) environments, in an ablation study where each of the algorithm’s components — centralised-$Q$ resampling (C-Q), adaptive training budget allocation (TR-A), and UCB-based validation budget allocation (VR-UCB) — is removed in turn.
Figure 11: Ablation of HYDRAVIPER on the physical deception (PD, left, higher rewards are better) and Ingolstadt corridor (IC, right, lower rewards are better) environments.